OpenStudy (anonymous):

what is the solution? solve for x 12 (3 - x) = 60

OpenStudy (anonymous):

Isolate x. Then Solve for it.

OpenStudy (rishabh.mission):

x=-2

OpenStudy (anonymous):

show the work?

OpenStudy (anonymous):

Divide by 12. Subtract 3 from both sides. and divide by -1

OpenStudy (amorfide):

let y=3-x 12y=60 get y on its own divide by 12 on both sides y=5 3-x=5 x=-2

OpenStudy (anonymous):

Really no need for substitution, but mathematically sound.

OpenStudy (amorfide):

sorry i have never done it via substitution just thought i would give it a go :P